# Directed and Undirected Vertex Connectivity Problems are Equivalent for Dense Graphs

### v1

@@ Abstract
Vertex connectivity and its variants are among the most fundamental problems in graph theory, with decades of extensive study and numerous algorithmic advances. The directed variants of vertex connectivity are usually solved by manually extending fast algorithms for undirected graphs, which has required considerable effort.
In this paper, we present a simple, black-box randomized reduction from directed to undirected vertex connectivity for dense graphs. As immediate corollaries, we largely simplify the proof for directed vertex connectivity in $n^{2+o(1)}$ time [LNP+25], and obtain a parallel vertex connectivity algorithm for directed graphs with $n^{\omega+o(1)}$ work and $n^{o(1)}$ depth, via the undirected vertex connectivity algorithm of [BJMY25]. Our reduction further extends to the weighted version of the problem. By combining our reduction with the recent subcubic-time algorithm for undirected weighted vertex cuts [CT25], we obtain the first subcubic-time algorithm for weighted directed vertex connectivity, improving upon a three-decade-old bound [HRG00] for dense graphs.
@@ 1 Introduction
Vertex connectivity and its variances are among the most fundamental graph problems and have received extensive studies over the last several decades [Kle69, Pod73, ET75, Eve75, Gal80, EH84, Mat87, BDD+82, LLW88, CT91, NI92, CR94, Hen97, HRG00, Gab06, CGK14, NSY19, FNY+20, SY22, NSY23, BJMY25, Kor25, CT25, JNSY25]. For an undirected graph, a vertex cut is a vertex set whose removal disconnects the graph. The problem of vertex connectivity asks to find a minimum vertex cut of a non-complete undirected graph. For directed graphs, a vertex cut is a vertex set whose removal renders the graph not strongly connected. The problem of directed vertex connectivity is to find a minimum vertex cut on a non-complete directed graph.
For a long time, progress in solving directed vertex connectivity often involved a manual, ad-hoc process of “mimicking” fast algorithms from the undirected graphs. For example, consider the fastest known algorithm [^1] Unless stated otherwise, all algorithms in this paper are randomized Monte Carlo algorithms. A gap remains between randomized and deterministic algorithms for vertex connectivity. We refer to [JNSY25] for more discussion. for vertex connectivity. A long line of research [NSY19, FNY+20, LNP+25] culminates in an almost linear-time randomized algorithm for undirected vertex connectivity. For directed graphs, [LNP+25] extends their techniques and spends 9 more pages to obtain an $n^{2+o(1)}$-time randomized algorithm for directed vertex connectivity, where $n$ denotes the number of vertices.
A similar phenomenon appeared in the algebraic approach, which was first proposed for undirected graphs [LLW88]; later, the same approach was manually adapted to work for directed graphs [CT91]. Very recently, this approach has been made faster on undirected graphs using common-neighborhood clustering techniques to obtain an $n^{\omega+o(1)}$-time algorithm [BJMY25] where $\omega\leq 2.371339$ is the matrix multiplication exponent [ADV+25]. It is conceivable that their techniques can also be extended to directed graphs, but this has not been done yet.
This suggested a deep connection between undirected and directed problems. In this paper, we resolve the missing link by introducing a simple black-box randomized reduction from directed to undirected vertex connectivity for dense graphs. This formalizes the long-observed phenomenon that the problems were related and explains why mimicking was so often successful.
@@
The directed vertex connectivity problem on an $n$-vertex directed graph reduces to undirected vertex connectivity problems with $O(n^{2}\log n)$ edges.
@@
As a corollary of Theorem 1.1, we obtain the following results immediately. Combining with $m^{1+o(1)}$ undirected vertex connectivity algorithm by [LNP+25], we obtain:
@@
Thus, it is no longer necessary to open the box and mimic the proof of undirected vertex connectivity as done in [LNP+25]. Furthermore, since the reduction is fully parallelizable, we also obtain parallel algorithms in this setting. Combining with $n^{\omega+o(1)}$-time undirected vertex connectivity algorithm by [BJMY25], we obtain:
@@
(Weighted) Vertex Mincut.
@@
Interestingly, our reduction Theorem 1.1 also holds on weighted graphs. Strictly speaking, we have the following result.
@@
The directed minimum vertex cut problem on an $n$-vertex directed graph reduces to the undirected minimum vertex cut problem with $O(n^{2}\log n)$ edges.
By plugging in Theorem 1.4 with the $\min(mn^{0.99+o(1)},m^{1.5+o(1)})$ undirected minimum vertex cut algorithm [CT25], we immediately get the following corollary that improves the $\widetilde{O}\left(mn\right)$ algorithm [HRG00] from 30 years ago on dense graphs.
@@
Concluding Remarks.
One can notice that our reduction from directed to undirected graph has a loss in boosting the number of edges to $n^{2}$, which is not a good reduction when the graph is moderately sparse. This also collides with the fact that directed vertex connectivity has remained at $n^{2+o(1)}$ running time for some time now. It is a great open problem to prove that directed vertex connectivity can be solved in almost linear time for any sparsity, not necessarily done by a black-box reduction. Another open problem is whether the reduction can be made deterministic. This seems plausible since vertex connectivity can be solved deterministically in $m^{1+o(1)}k$ time on unweighted undirected graphs where $k$ is the vertex connectivity, and $\tilde{O}(m^{1+o(1)}n)$ time on weighted directed graphs [JNSY25].
@@ 2 Preliminaries
Graphs.
In this paper, we consider both directed and undirected graphs $G=(V,E)$. $E\subseteq V\times V$ is the edge set, and when $G$ is an undirected graph, we consider $(u,v)$ and $(v,u)$ as the same edge. When we say $G$ is vertex-weighted, each vertex is assigned a weight $w_{G}(v)\in{\mathbb{Z}}^{+}$ (when $G$ is clear from the context, we omit $G$ and write $w(v)$). We always assume $w(v)$ is a polynomially bounded integer for simplicity. We write $G^{T}$ as the graph obtained by reversing the arc of $G$.
We use $N_{G}(u)\subseteq V$ to denote the neighborhood set of $u$ on $G$ when $G$ is an undirected graph. When $G$ is directed, we use $N^{out}_{G},N^{in}_{G}$ to denote the out and in neighbors. When we use $N^{out}_{G},N^{in}_{G}$ on an undirected graph, both of them refer to $N_{G}$. We write $N_{G}[u]=N_{G}(u)\cup\{u\}$, similar for other notations. When $G$ is clear from the context, we omit $G$ from the subscript. The minimum degree of a graph $G$ is defined as $\min_{v\in V}|N(v)|$, and the minimum weighted-degree is defined as $\min_{v\in V}(\sum_{u\in N(v)}w(u))$.
Vertex cuts.
A vertex cut $(L,S,R)$ is a tri-partition of the vertex set $V$ such that there are no edges from $L$ to $R$ where both $L$ and $R$ are non-empty sets. The vertex set $S$ is called a separator (following the literature, we also refer to it as a vertex cut). The size of the vertex cut is $|S|$ and the weight of the vertex cut is $w(S):=\sum_{v\in S}w(v)$. If $s\in L,t\in R$, then we call $(L,S,R)$ (or $S$) a $(s,t)$-vertex cut. It is an $s$-vertex cut if $s\in L$. Throughout this paper, we assume that every input graph is non-complete and thus a vertex cut always exists.
@@
The problem of single source minimum vertex cut is given a graph and a source $s\in V$, asks to find an $s$-vertex cut $(L,S,R)$ that minimizes $w(S)$ (in the case of a weighted graph) or the size $|S|$ (in the case of an unweighted graph). We denote $\kappa_{G}(s)=w(S)$ or $|S|$ in the case of an unweighted graph.
@@ 3 A Reduction from Directed to Undirected Vertex Connectivity
We state our main result.
@@
Suppose there is an algorithm ${\mathcal{A}}_{0}$ solving the minimum vertex cut on undirected graphs. In that case, there is a randomized algorithm solving minimum vertex cut on directed graphs w.h.p., making calls to ${\mathcal{A}}_{0}$ with cumulative $O(n^{2}\log n)$ many edges and additional $O(n^{2}\log n)$ time. Moreover, the reduction holds even when we restrict both inputs to unweighted graphs.
To prove Theorem 3.1, the reduction can be done in three steps as follows.
  1. 1.
    We reduce directed vertex connectivity (Directed VC) to directed single-source vertex connectivity (Directed SSVC) in Section 3.1.
  2. 2.
    Then, we reduce directed single-source vertex connectivity (Directed SSVC) to undirected single-source vertex connectivity (Undirected SSVC) in Section 3.2.
  3. 3.
    Finally, we reduce undirected single-source vertex connectivity (Undirected SSVC) to undirected vertex connectivity (Undirected VC) in Section 3.3.
The constructions in Sections 3.1 and 3.3 are folklore. The key novelty is the reduction in Section 3.2.
3.1 A Reduction from Directed VC to Directed SSVC
@@
If there is an algorithm ${\mathcal{A}}_{0}$ solving single-source minimum vertex cut on directed graphs, then there is a randomized algorithm solving minimum vertex cut on directed graphs w.h.p., making calls to ${\mathcal{A}}_{0}$ with cumulative inputs $O(m\log n)$ edges and additional $O(n^{2})$ time. Moreover, the reduction holds even when we restrict both input graphs to be unweighted or undirected.
@@
We start by describing the algorithm. We assume the graph is not a complete graph as otherwise there is no vertex cut. Let the minimum weighted-degree of the input graph be $\delta$, and let $W=\sum_{v\in V}w(v)$. The algorithm samples $O(\log n\cdot\frac{W}{W-\delta})$ many nodes independently, where each node $u$ get sampled with probability $w(u)/W$. For each sampled node $u$, define $G_{u}$ as the graph obtained from deleting every edge $(v_{1},v_{2})$ with $v_{1},v_{2}\in N_{G}(u)$ from $G$. The algorithm calls ${\mathcal{A}}_{0}$ on $G_{u}$ with source $u$ for every sampled node $u$, and returns the smallest vertex cut among all returns. We repeat the same algorithm for $G^{T}$ and take the smaller vertex cut returned.
Correctness.
Let $(L,S,R)$ be one of the smallest vertex cut. Suppose without loss of generality that $w(L)\geq w(R)$ (one of $G$ or $G^{T}$ must satisfy this condition). Notice that $\delta\geq w(S)$ because deleting vertices of weight $\delta$ disconnects the graph, and $S$ should be the smallest weight vertex set whose deletion disconnect the graph. Thus, we get $w(L)\geq(W-\delta)/2$. According to Chernoff’s bound, w.h.p. there is a sampled node $u\in L$, whose single-source minimum vertex cut has weight at most $w(S)$ since $S$ is clearly a single-source vertex cut in $G_{u}$ (since we only delete edges from $G$ to get $G_{u}$). To see that the returned cut in $G_{u}$ must be a cut in $G$, let $(L^{\prime},S^{\prime},R^{\prime})$ be a vertex cut in $G_{u}$ with $u\in L^{\prime}$. Then we have $N_{G}(u)\subseteq L^{\prime}\cup S^{\prime}$ since all edges adjacent to $u$ is present in $G_{u}$. Thus, we only delete edges between $L^{\prime}\cup S^{\prime}$ from $G$ to get $G_{u}$, which means $(L^{\prime},S^{\prime},R^{\prime})$ is still a cut in $G$.
Complexity.
We will analyze the expected size of the summation of all the graphs input to ${\mathcal{A}}_{0}$. Let $u$ be a sampled node. Fix an edge $e\in E$, let $I_{e}$ be the event that $e\in G_{u}$. Notice that if $I_{e}$ happens, then one of $e$’s end point must not be a neighbor of $u$. In other words, let $e=(v_{1},v_{2})$, then $u$ must be in either $V-N^{in}_{G}(v_{1})$ or $V-N^{in}_{G}(v_{2})$. Moreover, we have $w(V-N^{in}_{G}(v_{1})),w(V-N^{in}_{G}(v_{2}))\leq W-\delta$. Thus, the probability that $I_{u}$ happens is at most $2(W-\delta)/W$. Notice that the size of $G_{u}$ is at most $\sum_{e\in E}I_{e}$, which means the expectation of the size of $G_{u}$ can be upper bounded by $m\cdot 2(W-\delta)/W$. Recall that we sampled $O(\log n\cdot\frac{W}{W-\delta})$ many nodes, so the total number of edges among all $G_{u}$ has expectation $O(m\log n)$. Notice that building all the graphs $G_{u}$ can be done by brute force in $O(n^{2})$ time.
Notice that when $G$ is unweighted or undirected, the graph $G_{u}$ becomes unweighted or undirected as well. ∎
3.2 A Reduction from Directed SSVC to Undirected SSVC

### v2

@@ Abstract
Vertex connectivity and its variants are among the most fundamental problems in graph theory, with decades of extensive study and numerous algorithmic advances. The directed variants of vertex connectivity are usually solved by manually extending fast algorithms for undirected graphs, which has required considerable effort. In this paper, we present an extremely simple reduction from directed to undirected vertex connectivity for dense graphs. As immediate corollaries, we vastly simplify the proof for directed vertex connectivity in $n^{2+o(1)}$ time [LNP+25], and obtain a parallel vertex connectivity algorithm for directed graphs with $n^{\omega+o(1)}$ work and $n^{o(1)}$ depth, via the undirected vertex connectivity algorithm of [BJMY25].
Our reduction further extends to the weighted, all-pairs and Steiner versions of the problem. By combining our reduction with the recent subcubic-time algorithm for undirected weighted vertex cuts [CT25], we obtain a subcubic-time algorithm for weighted directed vertex connectivity, improving upon a three-decade-old bound [HRG00] for dense graphs. For the all-pairs version, by combining the conditional lower bounds on the all-pairs vertex connectivity problem for directed graphs [AGI+19], we obtain an alternate proof of the conditional lower bound for the all-pairs vertex connectivity problem on undirected graphs, vastly simplifying the proof by [HLSW23].
@@ 1 Introduction
Vertex connectivity and its variants are among the most fundamental graph problems and have received extensive studies over the last several decades [Kle69, Pod73, ET75, Eve75, Gal80, EH84, Mat87, BDD+82, LLW88, CT91, NI92, CR94, Hen97, HRG00, Gab06, CGK14, NSY19, FNY+20, SY22, NSY23, BJMY25, Kor25, CT25, JNSY25]. For an undirected graph, a vertex cut is a vertex set whose removal disconnects the graph. The problem of vertex connectivity asks to find a minimum vertex cut of a non-complete undirected graph. For directed graphs, a vertex cut is a vertex set whose removal renders the graph not strongly connected. The problem of directed vertex connectivity is to find a minimum vertex cut on a non-complete directed graph.
For a long time, progress in solving directed vertex connectivity often involved a manual, ad-hoc process of “mimicking” fast algorithms from the undirected graphs. For example, consider the fastest known algorithm for vertex connectivity. A long line of research [NSY19, FNY+20, LNP+25] culminates in an almost linear-time randomized algorithm for undirected vertex connectivity. For directed graphs, [LNP+25] extends their techniques and spends 9 more pages to obtain an $n^{2+o(1)}$-time randomized algorithm for directed vertex connectivity, where $n$ denotes the number of vertices.
A similar phenomenon appeared in the algebraic approach, which was first proposed for undirected graphs [LLW88]; later, the same approach was manually adapted to work for directed graphs [CT91]. Very recently, this approach has been made faster on undirected graphs using common-neighborhood clustering techniques to obtain an $n^{\omega+o(1)}$-time algorithm [BJMY25], where $\omega\leq 2.371339$ is the matrix multiplication exponent [ADV+25]. It is conceivable that their techniques can also be extended to directed graphs, but this has not been done yet.
This suggested a deep connection between undirected and directed problems. In this paper, we resolve the missing link by introducing a simple black-box reduction from directed to undirected vertex connectivity for dense graphs. This formalizes the long-observed phenomenon that the problems were related and explains why mimicking was so often successful.
@@
The directed vertex connectivity problem on an $n$-vertex directed graph reduces to the undirected vertex connectivity problem on a graph with $2n$ vertices and $O(n^{2})$ edges.
@@
As a corollary of Theorem 1.1, we obtain the following results immediately. Combining with the $m^{1+o(1)}$ undirected vertex connectivity algorithm by [LNP+25], we obtain:
@@
Thus, it is no longer necessary to open the box and mimic the proof of undirected vertex connectivity as done in [LNP+25]. Furthermore, since the reduction is fully parallelizable, we also obtain parallel algorithms in this setting. Combining with the $n^{\omega+o(1)}$-time undirected vertex connectivity algorithm by [BJMY25], we obtain:
@@
All-pairs minimum vertex cuts.
The simplicity of our reduction facilitates extending the results to many other settings. We believe that our reduction paves the way to showing many problems related to vertex cut are equivalent on dense undirected and directed graphs. The reduction simplifies another paper [HLSW23].
Theorem 1.4 (Informal version of Theorem 3.1).
The all-pairs minimum vertex cuts problem on an $n$-vertex directed graph reduces to the same problem on a $2n$-vertex undirected graph with $O(n^{2})$ edges.
In [AGI+19], they proved that all-pair minimum vertex cut has a conditional lower bound of $\widehat{\Omega}(n^{4})$ for combinatorial algorithms on directed graphs. One of the main contributions of [HLSW23] was to extend this lower bound to undirected graphs, which takes 7 pages of proofs. Our theorem Theorem 1.4 immediately implies the same result when combining with [AGI+19].
Weighted vertex mincuts.
@@
Our reduction Theorem 1.1 also holds on weighted graphs. Strictly speaking, we have the following result.
@@
The directed weighted minimum vertex cut problem on an $n$-vertex directed graph reduces to the undirected weighted minimum vertex cut problem on a graph with $2n$ vertices and $O(n^{2})$ edges.
By plugging in Theorem 1.5 with the $O(\min(mn^{0.99+o(1)},m^{1.5+o(1)}))$ undirected minimum vertex cut algorithm [CT25], we immediately get the following corollary that improves the $\widetilde{O}\left(mn\right)$ algorithm [HRG00] from 30 years ago on dense graphs.
@@
Remark.
Our construction extends to Steiner vertex cut problems as well, but this does not immediately simplify the construction of [HLSW23] because there was no known conditional lower bound for the directed Steiner vertex cut problem before their work.
@@
Independent Work.
Independently of our work, the running time of the new weighted vertex connectivity algorithms have been improved to $O\left(\min\{mn^{11/12+o(1)}\cdot d^{1/12},n^{2.677}\}\cdot(\log W)^{O(1)}\right)\leq O\left(mn^{0.976}\cdot(\log W)^{O(1)}\right)$, where $d$ is the average vertex degree in of the input directed graph and $W$ is the ratio between the largest and smallest vertex weight [CMT26]. They design algorithms that work on directed graphs, whereas our results are based on a black-box reduction to undirected graphs. Furthermore, they also show an algorithm for unweighted directed graphs with running time of $O(\min\{m^{1+o(1)}k,n^{2+o(1)}\})$ where $k$ is the vertex connectivity.
@@ 2 Preliminaries
In this paper, we consider both undirected and directed graphs $G=(V,E)$. For undirected graphs, we write the edge set $E\subseteq\binom{V}{2}$ as a set of two-element subsets of $V$ and we use $N_{G}(u)\subseteq V$ to denote the neighborhood set of $u\in V$ in $G$. For directed graphs, we write the edge set $E\subseteq V\times V$ as a set of ordered pairs of $V$ and we use $N^{\text{out}}_{G}(u),N^{\text{in}}_{G}(u)$ to denote the out- and in-neighbors of $u\in V$ in $G$. When $G$ is vertex-weighted, each vertex is assigned a weight $w_{G}(v)\in{\mathbb{Z}}^{+}$ (when $G$ is clear from the context, we omit $G$ and write $w(v)$). If each vertex is assigned weight $1$, then the weight of a vertex set is simply its cardinality.
Vertex cut.
A vertex cut $(L,S,R)$ is a tri-partition of the vertex set $V$ such that there are no edges from $L$ to $R$ and both $L$ and $R$ are non-empty sets. The vertex set $S$ is called a separator (following the literature, we also refer to it as a vertex cut). The size of the vertex cut is $|S|$ and the weight of the vertex cut is $w(S):=\sum_{v\in S}w(v)$. If $s\in L$ and $t\in R$, then we call $(L,S,R)$ (or $S$) a $(s,t)$-vertex cut. It is an $s$-vertex cut if $s\in L$. Throughout this paper, we assume that every input graph is non-complete and thus a vertex cut always exists.
@@
The problem of single source minimum vertex cut asks, given a graph and a source $s\in V$, to find an $s$-vertex cut $(L,S,R)$ that minimizes $w(S)$ (in the case of a weighted graph) or the size $|S|$ (in the case of an unweighted graph). We denote $\kappa_{G}(s)=w(S)$ or $|S|$ in the case of an unweighted graph.
All-pairs minimum vertex cut.
The problem of all-pairs minimum vertex cut asks, given a graph $G$, to find a minimum vertex cut $(L,S,R)$ such that $s\in L,t\in R$, for every pair of nodes $(s,t)\in V$ (or report that such cut does not exist). We denote $\kappa_{G}(s,t)=w(S)$ or $|S|$ in the case of an unweighted graph.
Steiner minimum vertex cut.
The problem of Steiner minimum vertex cut asks, given a graph $G$ and a terminal set $T\subseteq V$, to find a minimum vertex cut $(L,S,R)$ such that $T\cap L\neq\emptyset$ and $T\cap R\neq\emptyset$, that is, a minimizer of $\kappa_{G}(s,t)$ over all pairs $s,t\in T$.
@@ 3 A Reduction from Directed to Undirected Vertex Connectivity
We first state our main result and then describe the reduction.
@@
Given a vertex-weighted directed graph $G=(V,E,w)$ with $|V|=n$, there is a vertex-weighted undirected graph $G^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ with $2n$ vertices and $O(n^{2})$ edges such that the following directed vertex cut variants in $G^{\prime}$ can be solved using an algorithm for the undirected problem in $G$: global, all-pairs, single-source/sink, Steiner. For the Steiner variant, the terminal set $T^{\prime}$ given as input to the undirected problem has size $2|T|$, where $T$ is the terminal set given as input to the directed problem.
Construction.
Consider a vertex-weighted directed graph $G=(V,E,w)$. We define the vertex-weighted undirected graph $G^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ as follows. We create two copies $V_{\text{out}}$ and $V_{\text{in}}$ of $V$ and connect all vertices within each copy, creating two cliques of size $n$. We then add a perfect matching connecting each vertex with its copy. Finally, we add the directed edges in $G$ from copy $V_{\text{out}}$ to copy $V_{\text{in}}$. Formally, we have
  • $V^{\prime}=V_{\text{out}}\cup V_{\text{in}}$, where $V_{\text{out}}=\{v_{\text{out}}:v\in V\}$ and $V_{\text{in}}=\{v_{\text{in}}:v\in V\}$ are two copies of $V$;
  • $E^{\prime}=E_{\text{out}}\cup E_{\text{in}}\cup E_{\text{matching}}\cup E_{\text{out}\rightarrow\text{in}}$, where
    • $E_{\text{out}}=\binom{V_{\text{out}}}{2}$ and $E_{\text{in}}=\binom{V_{\text{in}}}{2}$ are the cliques over $V_{\text{out}}$ and $V_{\text{in}}$, respectively,
    • $E_{\text{matching}}=\{\{v_{\text{out}},v_{\text{in}}\}:v\in V\}$ is the perfect matching between $V_{\text{out}}$ and $V_{\text{in}}$, and
    • $E_{\text{out}\rightarrow\text{in}}=\{\{u_{\text{out}},v_{\text{in}}\}:(u,v)\in E\}$ are the directed edges from $G$ embedded in $G^{\prime}$;
  • $w^{\prime}(v_{\text{out}})=w^{\prime}(v_{\text{in}})=w(v)$ for all $v\in V$.
This construction should be contrasted with the standard split graph construction [DF56] where we split every vertex $v\in V$ to $v_{\text{out}}$ and $v_{\text{in}}$ and add appropriate edges, which was known since 1956; this construction has been used in many places including max-flow based algorithms (e.g., see [HRG00, Gab06, NSY19, FNY+20, LNP+25]). In other words, the graph $G^{\prime}$ is the split graph $SG$ after removing directions and adding cliques on $V_{\text{out}}$ and another one on $V_{\text{in}}$, see Figure 1 for illustration.
Figure 1: (Left) An input graph $G=(V,E,w)$. (Middle) The output graph $G^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$. (Right) The split graph construction (ignoring weights) SGSG given $G$. The split graph usually has the edge weight as capacity in one direction; our graph has vertex weights and no edge weights.
3.1 Basic Facts
We first make two simple observations about vertex cuts in $G^{\prime}$.
Proposition 3.2.
For every vertex cut $(L^{\prime},S^{\prime},R^{\prime})$ in $G^{\prime}$, we either have $L^{\prime}\subseteq V_{\text{out}},R^{\prime}\subseteq V_{\text{in}}$ or $L^{\prime}\subseteq V_{\text{in}},R^{\prime}\subseteq V_{\text{out}}$.
Proof.
There are no edges in $G^{\prime}$ between $L^{\prime}$ and $R^{\prime}$, so each clique must be a subset of $L^{\prime}\cup S^{\prime}$ or $R^{\prime}\cup S^{\prime}$. As $V_{\text{out}}$ and $V_{\text{in}}$ are cliques and form a partition of the vertex set $V^{\prime}$, the claim follows. ∎
Proposition 3.3.
Without loss of generality, we can write every vertex cut in $G^{\prime}$ as $(L^{\prime},S^{\prime},R^{\prime})$ such that $L^{\prime}\subseteq V_{\text{out}}$ and $R^{\prime}\subseteq V_{\text{in}}$. Moreover, if we put $L=\{v\in V:v_{\text{out}}\in L^{\prime}\}$ and $R=\{v\in V:v_{\text{in}}\in R^{\prime}\}$, then the sets $L$ and $R$ are nonempty and disjoint, and there is no edge $(u,v)\in E$ such that $u\in L$ and $v\in R$.
Proof.
The first part immediately follows from Proposition 3.2 and the fact that $G^{\prime}$ is undirected, so $L^{\prime}$ and $R^{\prime}$ can be swapped. To see the second part, assume for contradiction that there is some vertex $v\in L\cap R$, that is, $v_{\text{out}}\in L^{\prime}$ and $v_{\text{in}}\in R^{\prime}$. Since we have the perfect matching edge $\{v_{\text{out}},v_{\text{in}}\}\in E_{\text{id}}$ in $G^{\prime}$, this contradicts the fact that $(L^{\prime},S^{\prime},R^{\prime})$ is a cut. Now assume for contradiction that there is some edge $(u,v)\in E$ such that $u\in L$ and $v\in R$. Again, this contradicts the fact that $(L^{\prime},S^{\prime},R^{\prime})$ is a cut as there is the edge $\{u_{\text{out}},v_{\text{in}}\}\in E_{\text{out}\rightarrow\text{in}}$ in $G^{\prime}$. ∎
Next, we show the key property of our construction.
@@ 3.1 Basic Facts
Let $\emptyset\neq R_{\text{in}}\subseteq V_{\text{in}}$ and let $R=\{v\in V:v_{\text{in}}\in R_{\text{in}}\}$. We have $w^{\prime}(N_{G^{\prime}}(R_{\text{in}}))=w(N_{G}^{\text{in}}(R))+w(V)$.
@@
We proceed by splitting the neighborhood of $R_{\text{in}}$ into its parts within $V_{\text{in}}$, $R_{\text{out}}$ and $V_{\text{out}}\setminus R_{\text{out}}$, where $R_{\text{out}}=\{v_{\text{out}}:v_{\text{in}}\in R_{\text{in}}\}$. First, we consider $N_{G^{\prime}}(R_{\text{in}})\cap V_{\text{in}}$. As $V_{\text{in}}$ is a clique, we have $N_{G^{\prime}}(R_{\text{in}})\cap V_{\text{in}}=V_{\text{in}}\setminus R_{\text{in}}$ and hence $w^{\prime}(N_{G^{\prime}}(R_{\text{in}})\cap V_{\text{in}})=w^{\prime}(V_{\text{in}}\setminus R_{\text{in}})=w(V\setminus R)$. Second, we consider $N_{G^{\prime}}(R_{\text{in}})\cap R_{\text{out}}$. As the perfect matching edges $E_{\text{matching}}$ connect each $v_{\text{in}}\in R_{\text{in}}$ with its counterpart $v_{\text{out}}$, this set is exactly $R_{\text{out}}$, hence $w^{\prime}(N_{G^{\prime}}(R_{\text{in}})\cap R_{\text{out}})=w^{\prime}(R_{\text{out}})=w(R)$. Third, we consider $N_{G^{\prime}}(R_{\text{in}})\cap(V_{\text{out}}\setminus R_{\text{out}})$. These are neighours of $R_{\text{in}}$ from those edges in $E_{\text{out}\rightarrow\text{in}}$ that do not have an endpoint in $R_{\text{out}}$. More precisely, we have
$N_{G^{\prime}}(R_{\text{in}})\cap(V_{\text{out}}\setminus R_{\text{out}})$ $=\{u_{\text{out}}:\{u_{\text{out}},v_{\text{in}}\}\in E_{\text{out}\rightarrow\text{in}},\ u_{\text{out}}\in V_{\text{out}}\setminus R_{\text{out}},\ v_{\text{in}}\in R_{\text{in}}\}$
$=\{u_{\text{out}}:(u,v)\in E,\ u\in V\setminus R,\ v\in R\}$
$=N_{G}^{\text{in}}(R).$
It follows that $w^{\prime}(N_{G^{\prime}}(R_{\text{in}})\cap(V_{\text{out}}\setminus R_{\text{out}}))=w(N_{G}^{\text{in}}(R))$. Putting the three parts together, we conclude that
$w^{\prime}(N_{G^{\prime}}(R_{\text{in}}))=w(V\setminus R)+w(R)+w(N_{G}^{\text{in}}(R))=w(N_{G}^{\text{in}}(R))+w(V).\qed$
3.2 Global minimum vertex cut
The following lemma establishes a precise relationship between the minimum vertex cuts of $G^{\prime}$ and $G$.